OpenStudy (anonymous):

9 fourth root of 7 end root minus 8 cubed root of 7 end root plus 4 cubed root of 7 end root of 3 fourth root of 7

OpenStudy (anonymous):

http://gyazo.com/ea5b5ded06d846e02fc1a46c1f68b818 link to give visual

OpenStudy (anonymous):

I got C for this one, but I'm not sure

rishavraj (rishavraj):

u r right....

OpenStudy (anonymous):

thanks again rish!

rishavraj (rishavraj):

u r welcome...anytime!!!1

OpenStudy (anonymous):

i got D check for me please!!

rishavraj (rishavraj):

yup u r correct...

OpenStudy (anonymous):

I got B for this one

rishavraj (rishavraj):

u r wrong

OpenStudy (anonymous):

wait, is it supposed to be a positve 35?

rishavraj (rishavraj):

hey i didnt get it...what do u mean bu positive 35...where

OpenStudy (anonymous):

for the first part of the problem?

rishavraj (rishavraj):

what ur choice is....7*5 = 35

OpenStudy (anonymous):

Yeah ok. I got D then

rishavraj (rishavraj):

nope its not D

OpenStudy (anonymous):

ya idk then, i'll just skip it for now and go back to it in a bit

OpenStudy (anonymous):

I got B

rishavraj (rishavraj):

u r wrong try... (a + b)(a - b) = a^2 - b^2

rishavraj (rishavraj):

If u say then i can upload the pic of solution of previous question

OpenStudy (anonymous):

yes please show me the pic but okay i will try that right now

OpenStudy (anonymous):

so you got A??

rishavraj (rishavraj):

yeah...

OpenStudy (anonymous):

yeah i see where i was messing up looking at your work thanks! and for number 5 that i just sent you, I got D??

rishavraj (rishavraj):

yeah its D 54

OpenStudy (anonymous):

ok thank you !

rishavraj (rishavraj):

anyytime!!!!

OpenStudy (anonymous):

this is the last one i promise but i got D for this one too !

rishavraj (rishavraj):

u r wrong...and u can ask questions no need to promise....

OpenStudy (anonymous):

ok yeah i have no idea how to do this one

rishavraj (rishavraj):

wanna hav the pic??

OpenStudy (anonymous):

yeah that would help a lot

OpenStudy (anonymous):

ok yeah i see what you did

OpenStudy (anonymous):

i messed up all my parenthesis

rishavraj (rishavraj):

oo...

OpenStudy (anonymous):

yeah that helped me though, thanks !

rishavraj (rishavraj):

solved that system of linear equations question???

OpenStudy (anonymous):

Yeah I did :D

rishavraj (rishavraj):

ooo thats good...

OpenStudy (anonymous):

yeah, you helped me a lot ! I feel bad haha

rishavraj (rishavraj):

no need i enjoyed....heheehee

OpenStudy (anonymous):

:D thanks dude!

rishavraj (rishavraj):

anytime!!!! I told the same before

OpenStudy (anonymous):

i'll let you know if i have more questions!

rishavraj (rishavraj):

yup do let me know....

OpenStudy (anonymous):

D??

rishavraj (rishavraj):

yup u r correct

OpenStudy (anonymous):

also D?

rishavraj (rishavraj):

bingoo...u r right...hehehehe

OpenStudy (anonymous):

B??

rishavraj (rishavraj):

u r wrong...

rishavraj (rishavraj):

try it now...

OpenStudy (anonymous):

A?

rishavraj (rishavraj):

nooo....subtract correcly...

OpenStudy (anonymous):

i always get stuck doing these

rishavraj (rishavraj):

need any help???

OpenStudy (anonymous):

yeah i do please

rishavraj (rishavraj):

7/8 - 2/5 = (35 - 16)/40

OpenStudy (anonymous):

ok now i got D or x=19/40

rishavraj (rishavraj):

its not x = 19/40 its answer is x^(19/40)

OpenStudy (anonymous):

yeah thats what i meant sorry

OpenStudy (anonymous):

ok this is the last one i got A or D

rishavraj (rishavraj):

no need to be sorry......

rishavraj (rishavraj):

see a^-1 = 1/a and 1/(a)^-1 = a...

OpenStudy (anonymous):

So it is D?

OpenStudy (anonymous):

because -4*-4*-4*-4 is positive 256

rishavraj (rishavraj):

no thats not the reason....

OpenStudy (anonymous):

oh

rishavraj (rishavraj):

\[1/4^{-4} = 4^{4}\]

OpenStudy (anonymous):

O OKAY YEAH i forgot all about the first part ;-; i havent done this in forever

rishavraj (rishavraj):

no problem at all....

OpenStudy (anonymous):

well i am done for the night, i owe you man! thanks so much! you're a life saver!

rishavraj (rishavraj):

anytime!!!!again...good night..

OpenStudy (anonymous):

same to you! goodnight !

rishavraj (rishavraj):

dont wish me good night..its 8:40 am here right now

OpenStudy (anonymous):

ooo hahah well good morning to you!!

OpenStudy (anonymous):

and have a wonderful day!

rishavraj (rishavraj):

thanks and i will surely have the same!!!